(CO 5) The monthly utility bills in a city are normally distributed with a mean of $121 and a standard deviation of $41. Find the probability that a randomly selected utility bill is between $110 and $130.
Answer #1

Mean = 121

SD = 41

Z = (X - mean) / SD

P(110 < X <130) = P((110-121)/41 < Z < (130-121)/41)

= P(-0.27 < Z < 0.22)

= P(Z < 0.22) - P(Z < -0.27)

= 0.5871 - 0.3936

= 0.1935